OpenStudy (mikezack123):

Simplify: (4x − 6) + (3x + 6)

OpenStudy (mikezack123):

7x 7x − 12 x 7x + 12

OpenStudy (dls):

\[\Huge 4x − \cancel{6} + 3x + \cancel{6}=?\]

OpenStudy (mikezack123):

4x-3x

OpenStudy (mikezack123):

which is 1x

OpenStudy (mikezack123):

which is x

OpenStudy (mikezack123):

so its the 3rd answer?

OpenStudy (dls):

what?

OpenStudy (dls):

4x+3x=-1x? great.

OpenStudy (anonymous):

:)

OpenStudy (mikezack123):

? thats not one of the answers though...

OpenStudy (dls):

4+3 = 1? I mean seriously?

OpenStudy (anonymous):

the ans is 4x + 3x mate

OpenStudy (mikezack123):

lol you wrote it equals -1 in one of ur comments

OpenStudy (mikezack123):

you got me confused

OpenStudy (dls):

how can you get confused with simple addition.........

OpenStudy (mikezack123):

cuz ur an retriceole!

OpenStudy (mikezack123):

lol openstudy rewrites bad words as other words lol

OpenStudy (mikezack123):

but sholdint be 7x?

OpenStudy (mikezack123):

b-cuz 4+3= 7

OpenStudy (mikezack123):

@jim_thompson5910

OpenStudy (skullpatrol):

Yes.

OpenStudy (mikezack123):

yep, im super confused

OpenStudy (skullpatrol):

7x is the answer @MikeZack123

OpenStudy (mikezack123):

fer sure?

OpenStudy (anonymous):

the answer is 12x raise to 2 + 6x - 36

OpenStudy (mikezack123):

oh no

OpenStudy (mikezack123):

now im confused again

OpenStudy (skullpatrol):

Absolutely positive 7x is the answer

OpenStudy (skullpatrol):

Do you want a proof?

OpenStudy (mikezack123):

that'll make me happy... yes.

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

\((4x − 6) + (3x + 6) \) \( = 4x − 6 + 3x + 6 \) \( = 4x + 3x - 6 + 6 \) \(= 7x + 0 \) \(= 7x\)

OpenStudy (skullpatrol):